# Detectability of axion-like dark matter for different time-delay interferometry combinations in space-based gravitational wave detectors

### v1

@@ I Introduction
The nature of dark matter is one of the greatest mysteries in modern physics. The existence of dark matter has been confirmed by various observations, including galactic rotation curves [Zwicky:1933gu, Rubin:1970zza, Rubin:1980zd], the cosmic microwave background [Planck:2018vyg] and gravitational lensing [Walsh:1979nx, Clowe:2003tk, Markevitch:2003at]. However, we still do not know its microscopic properties, such as its mass, spin, and possible interactions with ordinary matter. Among various dark matter candidates, axion-like dark matter is one of the most compelling and widely studied theoretical candidates [Kim:2008hd, Marsh:2015xka].
@@
Multiple hypothetical interactions can be used to probe for axion-like dark matter, including its couplings with photons, gluons, spins, and gravitational waves [Gu__2024, kim2024oscillationsatomicenergylevels, gué2025searchqcdcoupledaxion, annurev:/content/journals/10.1146/annurev-nucl-102014-022120, Gavilan_Martin_2025, miller2025gravitationalwaveprobesparticle]. Here, we mainly consider the interaction between axion-like dark matter and photons. Based on this interaction, there are mainly two methods of detection. One of them is to utilize the conversion between photons and axion-like dark matter in the presence of a background magnetic field [PhysRevLett.51.1415], which forms the underlying principle of experiments such as CAST [CAST:2017uph]. Designed to search for X-rays produced by the conversion of solar axions, CAST has set stringent constraints on the parameter space of axion-photon coupling, and the next-generation experiment IAXO will push the sensitivity beyond current limits [Armengaud_2019]. Further detection efforts constitute a broad experimental portfolio, including ABRACADABRA-10 cm [PhysRevLett.122.121802, PhysRevLett.127.081801], “light shining through a wall” [PhysRevD.78.092003, PhysRevD.92.092002], and other initiatives [PhysRevD.42.1297, PhysRevLett.104.041301, 8395076]. Axion-like dark matter can also be generated in the extremely high-temperature environment within the interior of massive stars during their later stages of evolution, in which a small portion of axion-like dark matter flows out from the stellar atmosphere and decays into photons. These photons may be detected with terrestrial radio telescopes on Earth [PhysRevD.97.123001, PhysRevLett.121.241102, PhysRevD.99.123021, PhysRevD.101.123003, PhysRevD.102.023504].
@@
In addition, with the existence of the background axion field, photons with different chirality would exhibit different phase velocities due to the photon-axion coupling, which is called the axion-induced birefringence effect [sigl2019axionlikedarkmatterconstraints]. For linearly polarized light, this phase velocity difference is manifested as the rotation of the polarization plane when passing through the background field. The traditional astrophysical observation methods have been employed to search for the birefringence effect caused by axion-like dark matter, such as the study of cosmic microwave background (CMB) polarization [Fedderke_2019, HARARI199267, Fujita_2021] and astrophysical polarization measurements [Alighieri_2010, Liu_2020, Liu_2023, Chen_2022, Yuan__2021]. However, due to system uncertainties and astrophysical foregrounds, the sensitivities for such experiments are limited. Recently, new experimental schemes have been proposed, which utilize high-precision interferometric measurement of tiny phase shifts in optical cavities [PhysRevD.101.095034, PhysRevLett.102.202001, PhysRevLett.132.191002, pandey2024resultsaxiondarkmatterbirefringent, PhysRevD.100.023548] and laser interferometers [DeRocco:2018jwe, Nagano:2019rbw, michimura2025searchesultralightvectoraxion]. Among those detection methods, space-based laser interferometers have longer baselines compared to ground-based detectors, eliminating the noise caused by uncontrollable factors on the ground. Therefore, these interferometers may reach high sensitivities for axion-like dark matter at low frequencies.
Recent studies have explored the potential of space-based laser interferometers to detect axion-like dark matter [Yu:2024enm, Yao_2025, Gu__2025, yao2025axionlikedarkmattersearch]. These studies have adapted existing interferometers to make the detectors sensitive to the axion-induced birefringence effect and have assessed their corresponding sensitivities. However, there are still some gaps in the existing research on axion dark matter detection. For space-based laser interferometers, time delay interferometry (TDI) is essential to suppress dominant noise sources, including laser frequency noise and clock noise. While many different TDI combinations exist, their sensitivities for the detection of axion-like dark matter may vary across different frequency bands. Previous research has primarily focused on a limited set of TDI combinations, and the optimal TDI combination for specific frequency bands still needs to be studied. Here we systematically evaluate the sensitivities of the ASTROD-GW µHz-band interferometer for three extra TDI combinations: Monitor, Beacon, and Relay. Through a comparative analysis with LISA-like interferometers utilizing various TDI combinations, we find that Monitor and Beacon combinations provide markedly superior sensitivity at high frequencies. Specifically, they enhance the optimal sensitivity by about an order of magnitude and deliver this sensitivity across a wide frequency band.
@@
The paper is organized as follows. In Sec. II, we introduce the theoretical framework of axion-induced birefringence and the single-arm response signal of the detector. In Sec. III, we introduce the basic information of the ASTROD-GW mission and calculate the one-sided power spectral density of the signals for different TDI combinations. In Sec. IV, We calculate the sensitivity curves of ASTROD-GW, LISA, Taiji and TianQin, and conduct an comparative analysis around various interferometers and different TDI channels. Finally, we conclude in Sec. V. In this paper, we use natural units $c=\hbar=1$ and the metric signature as (,+,+,+)(-,+,+,+).
@@ II Axion-induced birefringence effect
In this section, we introduce the theoretical formalism of the axion-induced birefringence effect and derive the resulting phase shift for circularly polarized light. The Lagrangian describing the Chern-Simons interaction between the axion field and the photon field is given by [PhysRevD.41.1231, PhysRevD.43.3789]
@@
Since the arm length and the observation time of the interferometer in this study are much shorter than the coherence length and the coherence time of the axion, respectively, we can treat the axion field as a coherent field. Then the axion field can be approximately described by [Khmelnitsky_2014, Hui_2021, Yao:2024fie]
@@
$v_{\pm}=\frac{\omega}{k}\simeq 1\pm\frac{g_{a\gamma}\dot{a}}{2k}.$ (4)
@@
The phase shift caused by the birefringence effect is then given by [Gu__2025]
$\Delta\phi(t)=\frac{1}{2}g_{a\gamma}[a(t)-a(t-L)]\equiv\phi_{\rm single\ link}.$ (5)
@@
Such phase shift induced by the axion-induced birefringence effect can be detected by the interferometer. This provides a promising method to probe axion-like dark matter with laser interferometers [Yu:2024enm, Yao_2025, Gu__2025, yao2025axionlikedarkmattersearch].
@@ III.1 Space-based gravitational wave detectors
The basic plan of ASTROD was proposed in 1993 and has since remained in the conceptual, simulation, and laboratory research phases [Ni_1996, NI20031437]. In 1996, the Mini-ASTROD mission concept was proposed with the aim of testing general relativity and mapping the gravitational field of the solar system [Ni_1996ns]. In 2009, ASTROD-GW was proposed for the detection of gravitational waves. The arm length of this mission is designed as $2.6\times 10^{11}\rm m$, and the three spacecrafts operated near the Lagrange points L3, L4 and L5 of the Sun-Earth system, respectively [Ni_2024acg]. The schematic diagram of the detectors is depicted in Fig. 1.
When light propagates through the axion-like field, the birefringence effect gives rise to a periodic variation in its polarization angle. However, conventional interferometer designs are insensitive to such polarization shifts, which necessitates specific modifications. At present, there are two main modification schemes for laser interferometers. One approach entails embedding a dark matter interferometer capable of detecting rotations in the polarization plane [PhysRevLett.132.191002, yao2025axionlikedarkmattersearch]. Another method involves inserting waveplates into the interferometer, ensuring that both the transmitted and received laser beams are circularly polarized [PhysRevD.98.035021, Yao_2025, Gu__2025]. This work focuses on the second approach. In Fig. 2, we show the modified diagram of a single-link system. By introducing waveplates into the original design, both the transmitted and received light between spacecraft undergo conversion from linear to circular polarization.
@@
After suppressing laser frequency noise and clock frequency noise using TDI technique (detailed in Section III.2), the dominant noise sources limiting sensitivity are test mass acceleration noise and optical metrology system noise. For ASTROD-GW, the one-sided power spectral density of the test mass acceleration noise $S_{\rm acc}$ and the optical metrology system noise $S_{\rm oms}$ are given by [Wang:2023jct]
@@
Table 1: The specific design parameters of ASTROD-GW [NI_2013], LISA [amaroseoane2017laserinterferometerspaceantenna], Taiji [10.1093/ptep/ptaa083], and TianQin [TianQin:2015yph].
For LISA, Taiji, and TianQin, the power spectral densities of these noises are given by [babak2021lisasensitivitysnrcalculations]
@@ III.2 TDI combinations and signal responses
Time-delay interferometry (TDI) was proposed to address the problem that laser frequency noise in space-based gravitational wave detectors cannot be canceled directly. The core concept involves synthesizing a virtual equal-arm interferometer by time-shifting and linearly combining the measurement data from the unequal-arm links. This method allows for the suppression of laser frequency noise in post-processing while preserving the gravitational wave signal. Since its initial proposal by Tinto et al. [Armstrong_1999] in 1999, TDI has evolved through three generations over the past two decades. These correspond to the 1st-generation combinations for static constellations, the 1.5th-generation for rigidly rotating constellations [PhysRevD.69.022001], and the 2nd-generation that accounts for the relative velocities between spacecrafts due to orbital dynamics [PhysRevD.68.061303]. TDI continues to advance rapidly and can now effectively suppress not only laser frequency noise but also platform jitter and clock noise [Otto_2012, PhysRevD.98.042003]. Given the similarity in principle between detecting axion-like dark matter signals and gravitational waves with space interferometers, the TDI techniques developed for noise cancellation are directly applicable to axion-like dark matter detection.
@@
Currently, the most common combinations include Michelson, Sagnac, fully symmetric Sagnac, Beacon, Relay, and Monitor. The first three combinations have already been thoroughly studied in the application of detecting axion-like dark matter, and it has been found that the Sagnac combination can maximize the sensitivity of the interferometer for axion-like dark matter [Yao_2025, yao2025axionlikedarkmattersearch]. Here we mainly study the remaining three combinations and compare them with the Sagnac combination.
The TDI combinations Monitor, Beacon, and Relay consist of (E,F,G), (P,Q,R) and (U,V,W) [PhysRevD.62.042002, Tinto:2020fcc], respectively. Here we mainly consider E, P and U combinations, and other combinations can be obtained by permuting the indices . We present the space-time diagrams [Wang:2024ckw] for these TDI combinations in Figs. 35. Their expressions are given by [WANG2012211, PhysRevD.106.044054]
@@
For each TDI combination, the one-sided power spectral density (PSD) of the signal can be calculated as
$P_{s}^{C_{i}}(f)=2\frac{|\tilde{C_{i}}(f)|^{2}}{T}=\frac{\rho_{DM}g_{a\gamma}^{2}T}{\omega^{2}}|c_{i}|^{2},$ (17)
@@
On the other hand, the one-sided PSDs of the noises for these TDI combinations are given by [Yu:2023iog, Wang:2024hbh, PhysRevD.111.064056, PhysRevD.108.044075]
$P_{n}^{E}(f)$ $=P_{n}^{P}=16[2\sin^{4}(\pi fL)+\sin^{2}(2\pi fL)]S_{\rm acc}$
@@ IV Sensitivities to the axion-photon coupling
In this section, we define the theoretical expression for the detector sensitivity and present a comparison of the sensitivity curves for the four kinds of detectors under three TDI combinations. We also include the constraints for axion-photon coupling from CAST [2017] and SN1987A [Payez_2015] as a reference.
The signal-to-noise ratio (SNR) is defined as  [PhysRevX.4.021030]
@@
$g_{a\gamma}=\left(\frac{P_{n}^{C_{i}}\omega^{2}}{\rho_{DM}T|c_{i}|^{2}}\right)^{1/2}.$ (22)
@@
Figure 7: Sensitivity curves of various space-based laser interferometers to the axion-photon coupling $g_{a\gamma}$ under the Monitor-E combination. From left to right, they represent the sensitivities of ASTROD-GW (red), Taiji (orange), LISA (green), and TianQin (blue). The grey area and the yellow area represent the parameter space excluded by the CAST [2017] and SN1987A [Payez_2015] experiments, respectively.
Figure 8: Sensitivity curves of various space-based laser interferometers to the axion-photon coupling $g_{a\gamma}$ under the Relay-U combination.
@@
Since the arm length of ASTROD-GW is approximately 100 times longer than those of LISA, Taiji, and TianQin, its theoretical sensitivity is expected to improve by two orders of magnitude. However, as shown in Figs. 78, the sensitivity improvement of ASTROD-GW is limited to about two orders of magnitude in the frequency range of $10^{-7},10^{-3}]$ Hz. The reason is that, as frequency $f$ increases, the optical metrology system noise $S_{\rm oms}$ dominates. As shown in Fig. 6, in the frequency range of $10^{-3},10^{-1}]$ Hz, $S_{\rm oms}$ governs the sensitivity, while in the frequency range of $10^{-7},10^{-4}]$ Hz, the acceleration noise $S_{\rm acc}$ dominates. Due to the design of the ASTROD-GW mission, its optical measurement noise is approximately one to two orders of magnitude higher than that of other space-based detectors, while the test mass acceleration noise is about one order of magnitude lower than that of other detectors. This is the reason why ASTROD-GW has better sensitivity only in the low-frequency range.
V Conclusion

### v2

@@ I Introduction
The nature of dark matter is one of the greatest mysteries in modern physics. The existence of dark matter has been confirmed by various observations, including galactic rotation curves [98, 68, 67], the cosmic microwave background [3] and gravitational lensing [78, 17, 49]. However, we still do not know its microscopic properties, such as its mass, spin, and possible interactions with ordinary matter. Among various dark matter candidates, axion-like dark matter is one of the most compelling and widely studied theoretical candidates [65, 1, 21, 41, 50].
@@
Multiple hypothetical interactions can be used to probe for axion-like dark matter, including its couplings with photons, gluons, spins, and gravitational waves [30, 40, gué2025searchqcdcoupledaxion, 29, 28, 54]. Here, we mainly consider the interaction between axion-like dark matter and photons. Based on this interaction, there are mainly two methods of detection. One of them is to utilize the conversion between photons and axion-like dark matter in the presence of a background magnetic field [74], which forms the underlying principle of experiments such as CAST [6]. Designed to search for X-rays produced by the conversion of solar axions, CAST has set stringent constraints on the parameter space of axion-photon coupling, and the next-generation experiment IAXO will push the sensitivity beyond current limits [8]. Further detection efforts constitute a broad experimental portfolio, including ABRACADABRA-10 cm [62, 70], “light shining through a wall” [66, 11], and other initiatives [33, 9, 42]. Axion-like dark matter can also be generated in the extremely high-temperature environment within the interior of massive stars during their later stages of evolution, in which a small portion of axion-like dark matter flows out from the stellar atmosphere and decays into photons. These photons may be detected with terrestrial radio telescopes on Earth [37, 36, 69, 43, 12].
@@
In addition, with the existence of the background axion field, photons with different chirality would exhibit different phase velocities due to the photon-axion coupling, which is called the axion-induced birefringence effect [73]. For linearly polarized light, this phase velocity difference is manifested as the rotation of the polarization plane when passing through the background field. The traditional astrophysical observation methods have been employed to search for the birefringence effect caused by axion-like dark matter, such as the study of cosmic microwave background (CMB) polarization [24, 34, 27] and astrophysical polarization measurements [4, 46, 45, 16, 95]. However, due to system uncertainties and astrophysical foregrounds, the sensitivities for such experiments are limited. Recently, new experimental schemes have been proposed, which utilize high-precision interferometric measurement of tiny phase shifts in optical cavities [51, 52, 35, 63, 44] and laser interferometers [19, 55, 53]. Among those detection methods, space-based laser interferometers have longer baselines compared to ground-based detectors, eliminating the noise caused by uncontrollable factors on the ground. Therefore, these interferometers may reach high sensitivities for axion-like dark matter at low frequencies.
[…]
@@
The paper is organized as follows. In Sec. II, we introduce the theoretical framework of axion-induced birefringence and the single-arm response signal of the detector. In Sec. III, we introduce the basic information of the ASTROD-GW mission and calculate the one-sided power spectral density (PSD) of the signals for different TDI combinations. In Sec. IV, We calculate the sensitivity curves of ASTROD-GW, LISA, Taiji and TianQin, and conduct an comparative analysis around various interferometers and different TDI channels. Finally, we conclude in Sec. V. In this paper, we use natural units $c=\hbar=1$ and the metric signature as (,+,+,+)(-,+,+,+).
@@ II Axion-induced birefringence effect
In this section, we introduce the theoretical formalism of the axion-induced birefringence effect and derive the resulting phase shift for circularly polarized light. The Lagrangian describing the Chern-Simons interaction between the axion field and the photon field is given by [14, 15]
@@
Since the arm length and the observation time of the interferometer in this study are much shorter than the coherence length and the coherence time of the axion, respectively, we can treat the axion field as a coherent field. Then the axion field can be approximately described by [39, 38, 92]
@@
$v_{\pm}=\frac{\omega}{k}\simeq 1\pm\frac{g_{a\gamma}\dot{a}}{2k}=1\pm\delta v.$ (4)
@@
The phase shift caused by the birefringence effect is then given by [31]
$\Delta\phi(t)$ $=k\delta L(t)=k\int_{t-L}^{t}\delta vdt^{\prime}$
$=\frac{1}{2}g_{a\gamma}[a(t)-a(t-L)]\equiv\phi_{\rm single\ link}.$ (5)
@@
Such phase shift induced by the axion-induced birefringence effect can be detected by the interferometer. This provides a promising method to probe axion-like dark matter with laser interferometers [93, 91, 31, 89].
@@ III.1 Space-based gravitational wave detectors
The basic plan of ASTROD was proposed in 1993 and has since remained in the conceptual, simulation, and laboratory research phases [56, 57]. In 1996, the Mini-ASTROD mission concept was proposed with the aim of testing general relativity and mapping the gravitational field of the solar system [87]. In 2009, ASTROD-GW was proposed for the detection of gravitational waves. The arm length of this mission is designed as $2.6\times 10^{11}\rm m$, and the three spacecrafts operated near the Lagrange points L3, L4 and L5 of the Sun-Earth system, respectively [60]. The schematic diagram of the detectors is depicted in Fig. 1.
When light propagates through the axion-like field, the birefringence effect gives rise to a periodic variation in its polarization angle. However, conventional interferometer designs are insensitive to such polarization shifts, which necessitates specific modifications. At present, there are two main modification schemes for laser interferometers. One approach entails embedding a dark matter interferometer capable of detecting rotations in the polarization plane [35, 89]. Another method involves inserting waveplates into the interferometer, ensuring that both the transmitted and received laser beams are circularly polarized [20, 91, 31]. This work focuses on the second approach. In Fig. 2, we show the modified diagram of a single-link system. By introducing waveplates into the original design, both the transmitted and received light between spacecraft undergo conversion from linear to circular polarization.
@@
After suppressing laser frequency noise and clock frequency noise using TDI technique (detailed in Section III.2), the dominant noise sources limiting sensitivity are test mass acceleration noise and optical metrology system noise. For ASTROD-GW, the one-sided PSD of the test mass acceleration noise $S_{\rm acc}$ and the optical metrology system noise $S_{\rm oms}$ are given by [82]
@@
Table 1: The specific design parameters of ASTROD-GW [59, 83], LISA [5], Taiji [48], and TianQin [47].
For LISA, Taiji, and TianQin, the power spectral densities of these noises are given by [10]
@@ III.2 TDI combinations and signal responses
Time-delay interferometry (TDI) was proposed to address the problem that laser frequency noise in space-based gravitational wave detectors cannot be canceled directly. The core concept involves synthesizing a virtual equal-arm interferometer by time-shifting and linearly combining the measurement data from the unequal-arm links. This method allows for the suppression of laser frequency noise in post-processing while preserving the gravitational wave signal. Since its initial conceptual proposal in the 1990s [58, 57], TDI has evolved into three generations over the past two decades [22, 75]. These correspond to the 1st-generation combinations for static constellations, the 1.5th-generation for rigidly rotating constellations [72], and the 2nd-generation that accounts for the relative velocities between spacecrafts due to orbital dynamics [71]. TDI continues to advance rapidly and can now effectively suppress not only laser frequency noise but also platform jitter and clock noise [61, 77]. Given the similarity in principle between detecting axion-like dark matter signals and gravitational waves with space interferometers, the TDI techniques developed for noise cancellation are directly applicable to axion-like dark matter detection.
@@
Currently, the most common combinations include Michelson, Sagnac, fully symmetric Sagnac, Beacon, Relay, and Monitor. The first three combinations have already been thoroughly studied in the application of detecting axion-like dark matter, and it has been found that the Sagnac combination can maximize the sensitivity of the interferometer for axion-like dark matter [91, 89]. Here we mainly study the remaining three combinations and compare them with the Sagnac combination.
The TDI combinations Monitor, Beacon, and Relay consist of $\{E,F,G\}$, $\{P,Q,R\}$ and $\{U,V,W\}$ [23, 76], respectively. Each group of capital letters represents the signal combination number extracted from different spacecraft vertices in each TDI combination. Here we mainly consider $E$, $P$ and $U$ combinations, and other combinations can be obtained by permuting the indices. We present the space-time diagrams [83, 80, 84] for these TDI combinations in Figs. 35. Their expressions are given by [81, 79]
@@
For each TDI combination, the one-sided PSD of the signal can be calculated as
$P_{s}^{C_{i}}(f)=2\frac{|\tilde{C_{i}}(f)|^{2}}{T}=\frac{\rho_{\text{DM}}g_{a\gamma}^{2}T}{\omega^{2}}|c_{i}|^{2},$ (17)
@@
On the other hand, the one-sided PSDs of the noises for these TDI combinations are given by [94, 86, 97, 85]
$P_{n}^{E}(f)$ $=P_{n}^{P}(f)=16[2\sin^{4}(\pi fL)+\sin^{2}(2\pi fL)]S_{\rm acc}$
@@ IV Sensitivities to the axion-photon coupling
In this section, we define the theoretical expression for the detector sensitivity and present a comparison of the sensitivity curves for the four kinds of detectors under three TDI combinations. We also include the constraints for axion-photon coupling from CAST [7] and SN1987A [64] as a reference.
The signal-to-noise ratio (SNR) is defined as [13]
@@
$g_{a\gamma}=\left(\frac{P_{n}^{C_{i}}\omega^{2}}{\rho_{\text{DM}}T|c_{i}|^{2}}\right)^{1/2}.$ (22)
@@
Figure 7: Sensitivity curves of various space-based laser interferometers to the axion-photon coupling $g_{a\gamma}$ under the Monitor-E combination. From left to right, they represent the sensitivities of ASTROD-GW (red), Taiji (orange), LISA (green), and TianQin (blue). The grey area and the yellow area represent the parameter space excluded by the CAST [7] and SN1987A [64] experiments, respectively. The constraints from CMB polarization measurements are also shown in the plot. The purple line (BICEP/Keck results) represents the improved constraints on axion polarization oscillations in the CMB background. For detailed information, one can refer to [2]. Based on the Planck results [25], the cyan line and the pink dotted line represent the polarization washout effect and the Alternating Current (AC) oscillation effect with a rotation amplitude of at least $0.1^{\circ}$, respectively. The regions above the three diagonal lines mark the excluded parameter space for ultra-light axions.
Figure 8: Sensitivity curves of various space-based laser interferometers to the axion-photon coupling $g_{a\gamma}$ under the Relay-U combination. All other captions follow the same conventions as in Fig. 7.
@@
Since the arm length of ASTROD-GW is approximately 100 times longer than those of LISA, Taiji, and TianQin, its theoretical sensitivity is expected to improve by two orders of magnitude. However, as shown in Figs. 78, the sensitivity improvement of ASTROD-GW is limited to about two orders of magnitude in the frequency range of $10^{-7},10^{-3}]$ Hz. The reason is that, as frequency $f$ increases, the optical metrology system noise $S_{\rm oms}$ dominates. As shown in Fig. 6, in the frequency range of $10^{-3},10^{-1}]$ Hz, $S_{\rm oms}$ governs the sensitivity, while in the frequency range of $10^{-7},10^{-4}]$ Hz, the acceleration noise $S_{\rm acc}$ dominates. Due to the design of the ASTROD-GW mission, its optical measurement noise is approximately one to two orders of magnitude higher than that of other space-based detectors, while the test mass acceleration noise is about one order of magnitude lower than that of other detectors [59]. This is the reason why ASTROD-GW has better sensitivity only in the low-frequency range.
The coherence time $\tau_{c}$ and coherence length $\lambda_{c}$ of the axion-like dark matter are $\tau_{c}=\frac{2\pi\hbar}{mv_{\text{DM}}^{2}}\sim 1\text{yr}\left(\frac{10^{-16}\text{eV}}{m}\right),$ $\lambda_{c}=\frac{2\pi\hbar}{mv_{\text{DM}}}\sim 10^{4}\text{Gm}\left(\frac{10^{-16}\text{eV}}{m}\right)$, where $v_{\rm DM}\sim 10^{-3}$ represents the velocity of axion-like dark matter. When the coherence time and coherence length of the axion-like dark matter are much longer than the observation time $T$ and the detector arm length $L$, repectively, the spatial and temporal dependence of the axion-like dark matter can be neglected. For $T=1\,\text{yr}>\tau_{c}$, corresponding to axion-like dark matter with mass greater than $10^{-16}\,\rm eV$, the axion-like field cannot be regarded as a fully coherent field and it should be modeled as the superposition of several random harmonics [26]. One data processing method involves dividing the data set into several segments with durations less than $\tau_{c}$, searching for coherent axion dark matter signals within each of these segments. Then, a correction factor $(\frac{T}{\tau_{c}})^{\frac{1}{4}}$ needs to be included in the calculation of the sensitivity curve [92, 31, 13]. For $L>\lambda_{c}$, the axion-like field loses spatial coherence. In addition, due to the cancellation effect, the sensitivity curve will exhibit oscillations and have multiple singular points. Therefore, the sensitivity declines sharply in this region and needs to be corrected. We need to limit and truncate the detection range for each detector to ensure that the axion-like dark matter maintains coherence within its respective detection range.
V Conclusion and Discussion